OpenStudy (deadshot):

At the amusement park, you decide to ride the Ferris wheel, which has a maximum height of 50 meters and a diameter of 35 meters. It takes the wheel 11 minutes to make one revolution. Write the sinusoidal function, f(t), that models the height of your chair at any time, t.

OpenStudy (anonymous):

Sinusoidal function... \[f(t) = A \sin \omega t\]

OpenStudy (anonymous):

A can be determined from "maximum height of 50 meters and a diameter of 35 meters"... maximum height = 50 meter minimum height = 50-35 = 15 meter therefore what is A?

OpenStudy (anonymous):

One revolution = 11 minutes, this is equivalent to 1 period of a sine wave, from 50 meters then 15 meters then return to 50 meters again...

OpenStudy (anonymous):

from the information above, we can complete the sinusoidal function... :)

OpenStudy (anonymous):

... just the same as the simple harmonic motion we have previously (on your last problem)...

OpenStudy (deadshot):

so then it's \[f(t)=11\sin(\frac{ 2\pi }{ 35 })t+50\]?

OpenStudy (anonymous):

no... we said that for A, it is the amplitude of the function that varies from maximum to minimum... in your problem what quantity varies from maximum to minimum?

OpenStudy (deadshot):

the height of the chair

OpenStudy (deadshot):

so then would it be \[f(t)=17.5\sin(\frac{ 2\pi }{ 11 })t+15\] ?

OpenStudy (anonymous):



OpenStudy (deadshot):

The only options for amplitude on my homework are 11 and 17.5

OpenStudy (anonymous):

yes... that's correct... half of wheel diameter... and time t must be in minutes also....

OpenStudy (deadshot):

Alright, Thanks!